OpenStudy (anonymous):

some one study calculus 1

OpenStudy (chaise):

I can try help with your calculas problems - post a question.

OpenStudy (anonymous):

g(t) = t/2t+6 a) g(0) b) g(-3) c) g(10) d) g(x^2) e) g(t+h) f) g(t^2 - 3t +1)

OpenStudy (chaise):

This isn't calculus, this is basic manipulation of algebra. Simply make change your t values in the equation to what you have in the brackets. Eg: g(t) = t/2t+6 g(0) = 0/2x0+6 g(x^2) = x^2/2x^2+6

OpenStudy (anonymous):

sory